# Training-Free Token Pruning via Zeroth-Order Gradient Estimation in Vision-Language Models

### v1

@@ C.3 Evaluation Protocol and Benchmark Datasets
  • VQAv2-Test-Dev (Goyal et al., 2017): General-purpose VQA with real-world images and open-ended questions.
  • GQA (Hudson & Manning, 2019): Focused on compositional reasoning over scene graphs and structured images.
  • ScienceQA (IMG) (Lu et al., 2022): Multimodal science questions requiring domain knowledge and diagram interpretation.
  • TextVQA (Singh et al., 2019): Requires OCR capabilities to reason over text embedded within images.
@@
  • MMBench (Liu et al., 2024c): Assesses perception and reasoning across 20 fine-grained skill areas.
  • MMMU (Yue et al., 2024): Requires expert-level multimodal reasoning across 30+ subjects grouped into six major disciplines (e.g., Art & Design, Science, Engineering, Medicine), often involving complex diagrams and charts
  • SeedBench (Li et al., 2023a): Designed for evaluating multimodal large language models across diverse visually grounded question types, with an emphasis on perception, reasoning, and knowledge.
@@
To quantify the critical failure mode of object hallucination, we adopt the POPE (Li et al., 2023b) benchmark, which measures factuality in object recognition through binary existence questions. Performance is evaluated using the F1-score (F1) over object existence predictions, balancing precision and recall to reflect grounding reliability.
@@
For a holistic evaluation of both perceptual and cognitive abilities across numerous sub-tasks (e.g., OCR, counting, attribute recognition), we employ the MME (Zhang et al., 2021) benchmark. It reports separate scores for Perception and Cognition tasks, summed to form the combined Perception and Cognition score (P+C).
Figure A: FLOPs analysis of different token pruning methods on LLaVA-NeXT-7B across varying token retention rates. ZOO-Prune achieves similar FLOPs reduction to VisionZip and DivPrune while introducing only marginal overhead from sensitivity estimation.
@@
Figure B: Success cases illustrate that ZOO-Prune preserves key visual cues across attributes, spatial relations, and object identification, enabling accurate predictions even under aggressive pruning.
Figure C: Failure cases of ZOO-Prune, where predictions remain close in meaning to the reference answers yet are not exact matches. The examples illustrate typical situations where fine-grained distinctions, such as age or gender terms, lead to mismatches despite overall semantic proximity.
Appendix D Further Inference Efficiency Analysis
To further investigate the computational benefits of token pruning, we analyze the floating-point operations (FLOPs) required by different methods. Figure A reports the TFLOPs measured on LLaVA-NeXT-7B with varying token retention rates (640, 320, and 160). The baseline model without pruning requires 11.8 TFLOPs, while all pruning methods substantially reduce the cost. Specifically, VisionZip and DivPrune reduce the FLOPs to 2.61, 1.41, and 0.81 TFLOPs at 640, 320, and 160 tokens, respectively. Our ZOO-Prune achieves slightly higher FLOPs due to random perturbations introduced during sensitivity estimation, with 2.74, 1.53, and 0.94 TFLOPs, respectively. Nevertheless, the overhead is negligible compared to the baseline, and ZOO-Prune achieves comparable FLOPs reduction while consistently delivering stronger accuracy.
Appendix E More Visualization
E.1 More Qualitative Examples
Figure B provides extensive qualitative evidence for our method’s effectiveness across diverse examples, showcasing how its unified selection mechanism jointly leverages sensitivity and diversity to achieve precise token pruning. This mechanism is grounded in two concurrent principles. The zeroth-order sensitivity component identifies tokens that are most influential for answering a given question, ensuring the selection is rooted in semantic relevance. Simultaneously, the diversity criterion acts as a spatial and feature-space regularizer, guaranteeing the final token set is representative and non-redundant by preventing over-selection from visually similar regions, thereby preserving both local detail and global coverage.
The power of this dual-criterion approach is first demonstrated on questions that require fine-grained verification. For these tasks, such as confirming a person’s attributes or determining a relative spatial position like the catcher’s position, sensitivity is crucial for isolating the exact visual evidence. At the same time, diversity ensures that the context surrounding these key details is not entirely discarded, supporting a more robust and well-grounded verification, with correct predictions maintained even with a minimal number of retained tokens.
The method proves equally adept at handling open-ended identification tasks. Whether the challenge is to identify an object based on its relationship to others, like the car and laptop, or to recognize a subject in a scene, like the kite and deer, the fusion between the two criteria is vital. Sensitivity highlights the object of interest, while diversity provides a comprehensive, non-redundant set of its features, enabling the correct prediction. Collectively, the examples presented affirm that our sensitivity-aware diversity selection provides a principled way to maintain the visual context required for a wide spectrum of reasoning tasks.
E.2 Failure Case Examples
Figure C illustrates representative failure cases that, while penalized by the exact-match metric, simultaneously offer strong evidence for the effectiveness of our pruning strategy. The model’s predictions, such as “Baby” for “Boy” or “Woman” for “Lady,” are semantically correct and align with human interpretation, yet are deemed incorrect due to word-level differences. This exposes a limitation of evaluation protocols that reward surface-level string matching over semantic equivalence. Crucially, these predictions do not signal a failure in comprehension. Instead, they show that our sensitivity-aware pruning method successfully maintains the core semantics of the image even under aggressive compression. The model’s ability to generate plausible, near-synonymous concepts with only 64 tokens highlights the richness of the preserved representation and its support for high-level reasoning about key attributes like age and gender.
Appendix F The Use of Large Language Models (LLMs)
Throughout the preparation of this manuscript, we utilized Large Language Models (LLMs), specifically Google’s Gemini and OpenAI’s GPT-4, as writing assistance tools. Their primary role was to enhance the clarity, precision, and overall readability of the text. This included refining sentence structures, improving word choice, and ensuring grammatical correctness. In addition, the LLMs were occasionally employed to suggest alternative phrasings and to harmonize stylistic consistency across sections, thereby improving the presentation quality of the manuscript. All core research ideas, problem formulations, methodologies, experimental designs, results, and conclusions were entirely conceived, developed, and validated by the human authors. The LLMs were not involved in the generation of research content, analysis, or interpretation of findings. Instead, they served strictly as sophisticated language-polishing tools. The final version of the manuscript was carefully reviewed, edited, and approved by the authors to guarantee that it faithfully represents the original scientific contributions.

### v2

@@ C.3 Evaluation Protocol and Benchmark Datasets
  • VQAv2-Test-Dev [18]: General-purpose VQA with real-world images and open-ended questions.
  • GQA [19]: Focused on compositional reasoning over scene graphs and structured images.
  • ScienceQA (IMG) [28]: Multimodal science questions requiring domain knowledge and diagram interpretation.
  • TextVQA [36]: Requires OCR capabilities to reason over text embedded within images.
@@
  • MMBench [27]: Assesses perception and reasoning across 20 fine-grained skill areas.
  • MMMU [45]: Requires expert-level multimodal reasoning across 30+ subjects grouped into six major disciplines (e.g., Art & Design, Science, Engineering, Medicine), often involving complex diagrams and charts
  • SeedBench [20]: Designed for evaluating multimodal large language models across diverse visually grounded question types, with an emphasis on perception, reasoning, and knowledge.
@@
To quantify the critical failure mode of object hallucination, we adopt the POPE [22] benchmark, which measures factuality in object recognition through binary existence questions. Performance is evaluated using the F1-score (F1) over object existence predictions, balancing precision and recall to reflect grounding reliability.
@@
For a holistic evaluation of both perceptual and cognitive abilities across numerous sub-tasks (e.g., OCR, counting, attribute recognition), we employ the MME [50]. It reports separate scores for Perception and Cognition tasks, summed to form the combined Perception and Cognition score (P+C).
Table C: Performance Comparison on LLaVA-1.5-13B.
Method
GQA
Acc. $\uparrow$
MMB
Acc. $\uparrow$
MME
P+C $\uparrow$
POPE
F1 $\uparrow$
SQA
Acc. $\uparrow$
VQAV2
Acc. $\uparrow$
VQAText
Acc. $\uparrow$
MMMU
Acc. $\uparrow$
SEED-I
Acc. $\uparrow$
Avg.
$\uparrow$
Total 576 Tokens
Retain 192 Tokens $\downarrow$ 66.7%
Retain 128 Tokens $\downarrow$ 77.8%
Retain 64 Tokens $\downarrow$ 88.9%
(a) Prefilling Time (ms) vs. Performance (%) vs. TFLOPs
(b) E2E Latency (secs) vs. Performance (%) vs. TFLOPs
Figure A: Inference efficiency on the POPE benchmark relative to the LLaVA-NeXT-7B baseline. Circle size is proportional to TFLOPs. ZOO-Prune achieves the strongest accuracy–efficiency trade-off, sustaining higher POPE performance while operating at substantially lower latency and computational cost.
Appendix D Further Quantitative Results and Analysis
D.1 Additional Quantitative Results
Results on LLaVA-1.5-7B:
Table B reports the results on LLaVA-1.5-7B, the most widely used model in the LLaVA family. Across all token pruning levels, ZOO-Prune achieves the strongest overall performance. With a 66.7% pruning (192 tokens), it reaches an Avg. of 98.6%, surpassing the previous best method VisionZip (97.9%) and achieving leading scores on challenging tasks such as MMMU (36.11%) and SEED (58.80%). As pruning becomes more aggressive, the advantage of ZOO-Prune becomes even clearer. At 128 tokens, it maintains the best average score of 97.8%, substantially outperforming FastV (84.5%) and SparseVLM (93.0%). ZOO-Prune also delivers top results on SQA (59.49%), POPE (87.13%), VQAv2 (76.57%), and TextVQA (57.87%). Even under the extremely compressed 64-token setting, ZOO-Prune retains an average of 95.5%, outperforming the attention-based VisionZip by +2.4% and the diversity-based DivPrune by +0.7%. These results demonstrate that our sensitivity-aware diversity selection reliably preserves key visual cues and remains robust even under aggressive pruning.
Results on LLaVA-1.5-13B:
The robustness of ZOO-Prune is further validated in Table C, where methods are applied to the larger LLaVA-1.5-13B model. This evaluation tests the ability of pruning strategies to generalize to higher-capacity backbones. With 192 tokens retained, ZOO-Prune again obtains the highest average score (98.6%). Under the challenging 64-token setting, ZOO-Prune shows a clear performance advantage. It reaches 96.5% average accuracy, outperforming VisionZip by +2.8% and DivPrune by +1.1%. Retaining strong performance while discarding 88.9% of visual tokens highlights the robustness of our approach. Overall, across both 7B and 13B variants, ZOO-Prune consistently provides the best accuracy–efficiency trade-off, demonstrating strong generalization to different model capacities and pruning regimes.
Table D: Performance comparison between the projector and vision encoder layers as sensitivity proxies. The projector consistently provides the strongest accuracy across pruning ratios.
Vision Encoder
GQA
Acc. $\uparrow$
MMB
Acc. $\uparrow$
MME
P+C $\uparrow$
POPE
F1 $\uparrow$
SQA
Acc. $\uparrow$
VQAText
Acc. $\uparrow$
MMMU
Acc. $\uparrow$
SEED-I
Acc. $\uparrow$
Avg.
$\uparrow$
Total 2880 Tokens
LLaVA-NeXT-7B 64.20 67.90 1842.00 86.40 70.20 61.30 35.10 70.20 100%
Retain 640 Tokens $\downarrow$ 77.8%
Layer 01 61.23 65.29 1779.27 81.41 67.87 52.49 37.56 66.25 95.8%
Layer 03 61.96 65.64 1819.21 86.35 68.47 57.80 36.67 67.68 98.1%
Layer 05 61.73 65.55 1815.92 86.74 68.17 57.43 36.78 67.59 97.9%
Layer 07 61.85 65.29 1815.84 86.41 67.18 56.85 37.89 68.07 98.1%
Layer 09 62.18 65.12 1813.60 86.44 67.97 57.39 37.67 67.33 98.1%
Layer 11 61.98 65.21 1806.98 86.41 68.17 57.41 37.22 67.69 98.0%
Layer 13 61.93 65.12 1803.63 86.47 67.72 54.06 37.78 67.62 97.4%
Layer 15 62.03 65.38 1782.60 86.72 67.92 56.41 37.56 67.98 97.8%
Layer 17 61.99 65.38 1803.06 86.83 67.87 55.57 38.00 68.28 98.0%
Layer 19 61.94 65.03 1840.15 86.78 67.97 56.07 37.44 68.12 98.1%
Layer 21 62.03 65.29 1820.98 86.78 67.77 57.94 36.89 67.71 98.1%
Layer 23 61.70 65.46 1829.87 84.83 67.23 56.53 37.56 67.32 97.6%
Layer 24 61.99 65.21 1838.67 86.34 67.63 58.31 36.56 67.77 98.1%
Projector (Ours) 62.19 65.21 1816.45 86.75 68.02 57.98 36.89 67.95 98.2%
Retain 320 Tokens $\downarrow$ 88.9%
Layer 01 59.11 63.92 1664.19 76.08 68.86 49.38 36.89 63.94 92.4%
Layer 03 60.71 64.09 1796.69 84.08 67.87 56.74 36.44 65.64 96.3%
Layer 05 60.70 64.60 1789.64 84.48 67.82 55.77 38.44 65.77 96.9%
Layer 07 61.05 64.60 1746.22 83.23 67.67 55.36 37.11 65.87 96.0%
Layer 09 60.90 63.92 1798.27 83.74 67.53 55.81 37.00 65.62 96.2%
Layer 11 60.73 64.26 1784.26 84.36 65.80 56.13 36.56 65.80 95.9%
Layer 13 60.78 63.83 1712.90 83.76 67.38 52.13 36.67 65.64 94.7%
Layer 15 60.99 63.57 1754.49 84.44 67.53 55.02 37.22 66.39 96.0%
Layer 17 60.83 64.86 1799.67 85.00 67.33 53.70 37.11 66.48 96.3%
Layer 19 61.05 64.52 1801.79 85.44 67.63 54.51 37.67 66.42 96.8%
Layer 21 61.00 64.26 1764.22 84.37 66.93 56.05 37.67 65.72 96.4%
Layer 23 60.43 63.57 1766.28 81.46 68.12 54.95 36.89 65.19 95.3%
Layer 24 61.15 64.78 1775.97 84.23 67.03 56.14 37.00 66.29 96.4%
Projector (Ours) 60.97 64.86 1787.68 85.47 67.77 57.28 37.00 66.47 97.1%
Retain 160 Tokens $\downarrow$ 94.4%
Layer 01 57.62 61.34 1590.15 70.07 68.12 46.57 36.67 61.45 89.1%
Layer 03 58.88 63.14 1689.52 79.73 67.13 55.26 35.67 62.99 93.2%
Layer 05 59.69 62.54 1705.96 80.85 68.07 54.47 36.11 63.99 93.9%
Layer 07 59.43 63.92 1715.39 78.50 67.38 52.76 39.00 63.66 94.3%
Layer 09 59.58 62.46 1712.23 79.35 67.63 53.39 35.67 63.42 93.1%
Layer 11 59.18 63.23 1700.83 80.18 63.59 54.22 37.22 63.59 93.3%
Layer 13 59.33 62.80 1690.78 79.55 67.77 49.61 38.00 63.60 93.1%
Layer 15 59.78 63.14 1685.21 82.00 66.98 52.99 36.44 64.51 93.8%
Layer 17 60.06 63.32 1712.95 82.54 66.68 50.78 37.44 64.78 94.0%
Layer 19 59.90 62.63 1724.21 82.65 67.63 51.40 37.56 64.43 94.2%
Layer 21 59.53 63.66 1720.90 80.69 67.72 53.94 36.78 63.46 94.1%
Layer 23 58.80 62.71 1665.75 75.96 68.17 52.13 36.67 62.43 92.2%
Layer 24 59.57 63.14 1714.08 81.01 67.67 55.00 37.00 63.87 94.4%
Projector (Ours) 59.93 64.18 1738.64 83.05 68.42 55.42 37.11 64.05 95.4%
D.2 Further Inference Efficiency Analysis
[…]
Figure B: Revisiting Token Sensitivity Proxy: Projector vs. Encoder. Compared to the visual encoder, the projector is a reliable proxy for zoo-based token sensitivity estimation.
D.3 Visual Sensitivity: Projector vs. Encoder
The projection layer acts as a modality-alignment bottleneck, where high-level visual features are consolidated and mapped into the LLM embedding space. Tokens that remain influential at this stage are therefore those most critical for downstream language generation. Motivated by this, ZOO-Prune computes ZOO-based sensitivity at the projector rather than across the full vision encoder, avoiding costly end-to-end perturbations.
To evaluate whether the projector serves as a reliable proxy, we compare performance when sensitivity is computed at each encoder layer. As shown in Table D, mid-to-deep encoder layers (e.g., 17–24 layers) provide moderately strong estimates but consistently underperform the projector, while early layers degrade sharply under aggressive pruning (e.g., Layer 1 falls to 89.1% at a 160-token budget). Fig. B further highlights this trend: the projector achieves the highest and most stable accuracy across all pruning levels, with its advantage growing more pronounced as token budgets shrink. Overall, these findings show that the projection layer delivers the most stable, accurate, and efficient signal for token sensitivity, motivating its use as the default estimation strategy in ZOO-Prune.
Figure C: Selected visual tokens comparison on the GQA benchmark. Token pruning driven by text-visual (T2V) attention often suffers from positional bias, while visual-visual (V2V) attention tends to retain redundant token clusters. Diversity-based pruning spreads tokens broadly but lacks semantic focus. ZOO-based sensitivity can capture output-related tokens but overlooks spatial coverage. Our ZOO-Prune jointly optimizes sensitivity and diversity for balanced selection across compression ratios.
Figure D: Qualitative comparison of pruned token masks from VisionZip, DivPrune and ZOO-Prune on the GQA benchmark.
Figure E: Failure cases of ZOO-Prune illustrating two typical error patterns. Left: Predictions remain semantically close to references but differ in fine-grained age distinctions (e.g., “boy” vs. “baby”). Right: In visually cluttered scenes with multiple objects, the model makes incorrect predictions due to confusion among different items present in the scene (e.g., “mug” vs. “glass”).
@@ D.2 Further Inference Efficiency Analysis
Figure F: Additional success examples demonstrating that ZOO-Prune preserves key visual cues across attributes, spatial relations, and object identification, enabling accurate predictions even under aggressive pruning.
Appendix E Additional Qualitative Results
Sensitivity vs. Attention vs. Diversity.
Fig. C compares token selection patterns driven by different token importance strategies. Attention-based methods exhibit clear limitations: text-visual (T2V) attention often suffers from positional bias (e.g., disproportionately focusing on lower image regions, leading to the omission of key contextual cues), while visual-visual (V2V) attention tends to form redundant clusters. In contrast, diversity-based pruning treats all tokens equally, lacking semantic focus, whereas sensitivity-based selection captures key semantics but may result in spatial clustering. ZOO-Prune addresses these shortcomings by unifying sensitivity and diversity. This ensures that selected tokens are both semantically informative and spatially distributed, maintaining robust coverage of task-relevant regions even under an aggressive 64-token budget.
More Visualization Examples.
Fig. D visualizes the pruning masks and QA results. VisionZip prioritizes visual saliency, often focusing on the high-contrast kitchen background while missing the foreground “Bench” across all ratios. DivPrune enforces diversity but compromises object integrity (e.g., it sparsifies “Dog”-related tokens, leading to a Bear” hallucination). ZOO-Prune overcomes these limitations by using sensitivity to anchor dense token clusters on semantic targets. It effectively preserves the visual cues of key objects, ensuring correct predictions even when 88% of tokens are removed. This localization capability is further corroborated by Fig. F, demonstrating that ZOO-Prune consistently retains semantically critical tokens.
Failure Case Examples.
Fig. E illustrates representative failure cases under aggressive pruning, where ZOO-Prune produces semantically close but inexact predictions. These errors reveal two limitations: difficulty in capturing fine-grained attributes for distinguishing closely related concepts, and vulnerability to visual clutter with multiple interfering objects. While our method preserves high-level semantic integrity at 64 tokens, extreme reduction inevitably compromises nuanced visual cues required for precise recognition in complex scenes.
Appendix F Discussion and Future Work
Although ZOO-Prune achieves strong gains across vision–language reasoning tasks, several avenues remain open. First, our evaluation focuses on encoder–decoder VLMs such as LLaVA-NeXT. Extending ZOO-Prune to broader architectures, including emerging Omni-style unified models, may reveal different sensitivity patterns and grounding behaviors. Second, the method is currently image-centric. Applying ZOO-Prune to video, 3D scenes, or egocentric data will require reconsidering token selection under temporal and geometric structure. Finally, integrating ZOO-Prune into Vision-Language-Action agents is a promising direction, as adaptive preservation of task-critical visual cues may improve long-horizon stability and reduce cascading errors. In sum, extending ZOO-Prune toward Omni-style models, richer modalities, and interactive agents presents a promising direction for broadening the scope and impact of test-time visual refocusing.